40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6065     Entry time: Sat Dec 3 18:29:20 2011
Author: Den 
Type: Update 
Category: Adaptive Filtering 
Subject: coherence 

I've looked through the coherence between the MC length and seismometers after the if-statement problem was fixed. Coherence improved for all seismometers but is still not 1. It is possible that contribution from X, Y, Z directions split the coherence between them but at ~0.2-03 Hz we do not see much coherence for all these directions.

coh_mcl_seis.pdf

I looked at the coherence between MC2 OSEM signal and MC_F when the AUTO LOCKER is ON and OFF. I thought that we'll ses the same coherence for both regimes as laser is locker to the MC length. However, I figured out the coherence is worse when AUTO LOCKER is ON at frequencies 0.2-0.3 Hz.

sensor_locker.pdf

The first idea that comes to mind is that when feedback to the laser is provided, the pressure to the mirrors from the laser beam is changed.

ELOG V3.1.3-